# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v3

@@ V-A2 Hypothesis
We expect PARC to be able to compute plans to the goal reliably, with no safety violations during $[0,t_{\mathrm{\textnormal{f}}}]$. However, due to the naïve choice of planning model, we expect the avoid set to be overly conservative, especially since (54) has three non-ETI states ($v_{x},v_{z},\omega$) in its PWA formulation.
@@ V-A4 Discussion
In Appendix -G and the next few examples, we show how a careful design of the planning model leads to much better performance on more complicated scenarios, as well as enabling stabilizing behavior such as braking without explicitly modeling velocity as a goal state.
@@ V-B Impact of Planner-Tracker Cooperation
We now study cooperative vs. adversarial planning and tracking models. We also further study the Neural CLBF [dawson2022safe] learning-based reach-avoid approach. For this experiment, we use a 3D near-hover quadrotor introduced for FaSTrack [chen2021fastrack], a reachability-based planner tracker that uses adversarial planning and tracking models.
@@ V-B1 Experiment Setup
The tracking model and PARC’s planning model are listed in Appendix -E. The quadrotor must navigate a narrow gap (width 1.8 m, compared to drone width 0.54 m) to reach its goal as shown in Fig. 7. We simulate 8,100 trials from different initial conditions. We measure the number of trials that reach the goal, and the number that crashes. We also consider the total computation time of each method. Further experiment details are in Appendix -F.
@@ V-B2 Hypotheses
We anticipate that our PARC planner always reach the goal safely. We also anticipate faster computation time of our reachability analysis versus the other two approaches, but note that FaSTrack does not need to recompute the reachable set for every obstacle. Since FaSTrack provides formal proof of safety, we anticipate it has no crashes. However, since it treats the planner and tracker as adversarial, we expect it to struggle to reach the goal. On the other hand, since CLBFs specifically balance safety and liveness, we expect it to reach the goal more often than FaSTrack but at the expense of crashing due to the high chance of learning an incorrect representation of safety when forced to operate near obstacles.
@@ V-B3 Results
Fig. 8 illustrates the trajectories simulated on the grid of 8,100 initial positions using each method. Neural CLBF found 576 successful trajectories (7.1 %), while safety was violated for the rest. FaSTrack found 0 successful trajectories but never violated safety. However, PARC found the plan to reach the goal without collision for 2,592 initial states (32 %).
@@
Figure 8: Simulated trajectories of 10-D quadrotor for 8,100 initial states using three methods: PARC, FaSTrack [chen2021fastrack], Neural CLBF [dawson2022safe]. The green box is the goal set and the pink boxes are obstacles.
@@ V-B4 Discussion
The performance of FaSTrack is notably compromised by the large, conservatively computed TEB, specified as a $2.47\times 2.47\times 0.1$ box. This often results in the inability to compute feasible plans in spaces tighter than the TEB. However, as seen in Fig. 9, sampled tracking error under parameterized plan, the tracking error is much smaller than TEB, enabling PARC to find a plan in challenging scenarios.
Although Neural CLBF succeeds in capturing most of the safe, unsafe, goal regions, the synthesized control policy steers some of the initial states to violate safety. Thus, the learning-based method sacrifices guarantees.
PARC has a strong safety guarantee, so every plan PARC found did not collide with the obstacles. We note that a better choice of planning model could increase PARC’s success rate; the simplicity of the single-integrator planning model, employed mainly for fair comparison with FaSTrack, reduced the flexibility of possible plans. Indeed we leverage a better choice of trajectory model next.
TABLE II: Performance Evaluation of PARC and Other Reach-Avoid Frameworks: This table compares success and safety rates across various systems and scenarios. The success rate is calculated as the percentage of simulations achieving their goals without collisions, while the safety rate measures the avoidance of obstacles. Computation times are categorized into offline and online phases, with the offline phase not involving known goals, obstacles, or initial states. We report the mean and the standard deviation for time.
@@
13D Quadrotor PARC-Funnel 1.00 1.00 $3.07\pm 0.06$ 00 V-D
@@ V-C1 Experiment Setup
We use the same planning model as quadrotor RTD [kousik2019safe], which generates time-varying polynomial positions in 3-D [mueller2015computationally] with $\Delta t=0.02$units with $t_{\mathrm{\textnormal{f}}}=3$units; this formulation is affine time-varying in the trajectory parameters with no non-ETI states, so reachability can be tightly approximated by PARC. Both PARC and RTD use the same tracking error that is precomputed in [kousik2019safe].
@@
We declare safe if the quadrotor does not crash, and success if the quadrotor reaches the goal region without colliding with any obstacles.
@@ V-C2 Hypothesis
We anticipate that PARC will always safely find a way to reach the goal since we know the narrow gap is large enough to do so, and PARC’s primary source of conservativeness is its tracking error estimation (which is identical to that of RTD). We anticipate that RTD will reach the goal much less often because it suffers an additional course of conservativeness from using a polynomial planning model with zonotopic reachable sets.
@@ V-D Impact of Time-Varying Tracking Error
Figure 11: Comparison of PARC (a-b) and KDF (c-d) motion planning for a 3-D quadrotor across two scenarios: a narrow gap of $0.85m$ (left) and a wide gap of $3m$ (right), projected onto the $p_{x}$-$p_{y}$ plane. The initial planning state is $\mathbf{p}_{0}=[4.8,1,4]^{{}^{\top}}$ with a planning horizon of $t_{\mathrm{\textnormal{f}}}=15$units. Goals are in the green box, the dense pink boxes indicate the actual obstacles and the lighter pink boxes show the obstacles buffered by the drone’s volume. Pink trajectories within blue volumes represent computed motion plans and bounds of realized trajectories when tracking the plan (i.e. funnels), illustrating how each method handles tracking errors at the motion-planning level.
Finally, we explore the importance of accounting for the temporal correlation of tracking errors. In this experiment, we compare our approach with the KDF framework that combines sampling-based planning with funnel-based feedback control [verginis2022kdf]. This control mechanism prescribes that tracking error remain in funnels described by a performance function, $\rho(t)$, which decays exponentially over time [bechlioulis2014low]. Employing the same funnel-based control across both the KDF and PARC methodologies allows us to maintain a consistent tracking error model. This setup confirms that any observed differences in performance are solely attributed to the planner’s strategy of using tracking errors.
Key result: By considering time-varying as opposed to worst-case tracking error, PARC significantly reduces conservativeness of motion planning. Furthermore, PARC’s use of parameterized trajectories enabled the design of a funnel controller that maintained the prescribed tracking error bounds, whereas the same controller could not reliably track the smoothed output of a sampling-based motion planner.
@@ V-D1 Experiment Setup
We extend the setup from Section V-B to compare PARC and KDF using gap widths of 0.85 m, and 3 m, relative to the quadrotor’s 0.54 m of body width. We evaluate success and safety ratios, along with computation times, by simulating the quadrotor from zero velocity at 3,700 and 2,400 uniformly distributed initial positions across each scenario.
@@
For the tracker, both PARC and KDF use a specialized funnel controller adapted for an under-actuated 3-D quadrotor proposed in [lapandic2022robust], as the conventional KDF funnel-based feedback controller is unsuitable for such systems [lapandic2023kinodynamic]. This controller conforms to the performance function $\rho(t)=\rho_{0}+(\rho_{0}-\rho_{\infty})e^{-\lambda t}$ where we choose $\rho_{0}=0.5,\rho_{\infty}=0.1,\lambda=0.5$. We note that this controller does not respect input limits. The 48 hyperparameters for the funnel controllers are reported in Appendix -F.
Despite extensive tuning efforts, a universal hyperparameter set for KDF could not be determined that respected the prescribed tracking error bounds; thus, we assume the ideal performance of the funnel controller, and base our statistics on reference trajectories ignoring tracking error. In other words, the comparison is biased towards benefitting KDF and disadvantaging PARC. We note that we designed controllers that function for a selected few initial states with KDF, the outcomes of which are demonstrated in Appendix -F and Fig. 11. In contrast, since PARC uses parameterized reference trajectories, we were able to succesfully identify a universal set of hyperparameters that enabled us to base the statistics on realized trajectories that included tracking error.
@@ V-D2 Hypothesis
In the wide-gap scenario, both PARC and KDF are hypothesized to successfully execute reach-avoid plans. We anticipate KDF to demonstrate faster computation times, benefiting from its sampling-based planning approach. In the narrow-gap scenario, KDF is expected to struggle due to its reliance on maximum funnel values for free space construction in RRT, whereas PARC is predicted to succeed by leveraging the time-varying nature of tracking errors. We expect no safety violations for all experiments.
@@ V-D4 Discussion
As anticipated, PARC demonstrates superior performance over KDF in narrow-gap scenarios due to its ability to utilize the time-varying nature of tracking errors effectively; thus taking advantage of a low tracking-error regime. This advantage is visually depicted in Fig. 11 a, where PARC waits for the funnel controller to reduce the tracking error before navigating through narrow gaps. In contrast, for wider gaps shown in Fig. 11 b, the quadrotor proceeds immediately. Conversely, KDF struggles in narrow spaces, as depicted in Fig. 11 c, because its planner computes paths based on a static maximum tracking error that exceeds the actual gap width (i.e., $\rho_{0}=0.5>0.31$), but performs well in wider spaces where this limitation does not impede planning as depicted in Fig. 11 d.
@@
PARC achieves reach-avoid plans in narrow-gap scenarios but faces lengthy computation times due to detailed sampling from BRAS and smaller $\Delta t$ steps. In contrast, KDF uses sampling-based motion planning (SBMP) effectively in scenarios with adequate gap widths, swiftly creating viable plans. This indicates a potential research direction: integrating SBMP within PARC could shorten computation times and facilitate the generation of expert trajectories that adeptly handle near-danger scenarios, surpassing conventional kinodynamic SBMP methods such as KDF.
@@ V-E Overall Discussion
Through the above experiments, we have shown how our proposed PARC approach confers advantages in computing and using the Backward Reach-Avoid Set (BRAS). In near-danger scenarios, PARC was less conservative than FaSTrack and KDF since the former uses a time-variant tracking error function, whereas the latter two assumes constant, worst-case tracking errors. While RTD can similarly handle time-variant tracking error functions, PARC still outperforms it due to the lower numerical approximation error from representing the planning model as a PWA system instead of a polynomial and using H-polytopes instead of zonotopes. Finally, Neural CLBF cannot provide hard reach-avoid guarantees, resulting in many crashes with obstacles.
However, PARC still has room for improvement. Most obviously, the safe plan computation time is on the order of seconds for just a few obstacles; this is too slow for real-time operation. In contrast, RTD [kousik2019safe], although more conservative and failed to reach the goal, can compute a new plan about every 51 ms in Section V-C. Typically, we expect motion planners with real-time capabilities to compute new plans for at least every 0.5 s [kousik2020bridging]. To achieve this for PARC, parallelization needs to be implemented as a future work. Furthermore, the planning model requires careful hand-crafting as shown in our experiments.
@@ VI-A2 Reach-Avoid Method Comparison
Our drifting tracking model dynamics (see (96) below) are not control affine, so directly applying a CBF or CLBF control approach is not tractable [ames2016control, dawson2023safe]; future work may apply results for non-control-affine systems [son2019safety, seo2022safety]. Note, [dawson2022safe] poses a control affine vehicle dynamics model but does not include tire saturation, so cannot drift. Furthermore, our planning and tracking models are too high-dimensional for standard HJB value function computation [chen2021fastrack, mitchell2008flexible]; future work will compare against learning-based HJB [bansal2021deepreach]. The most similar work is RTD [kousik2020bridging, liu2022refine], which has not been applied to these extreme dynamics. As a future implementation, RTD’s reachable set can use PARC’s formulation to enable a direct drifting comparison; then, the only difference with PARC is that RTD implicitly represents the avoid set. Due to this similarity, we focus on demonstrating PARC’s capabilities as opposed to a head-to-head comparison.
@@ VI-B Demo Details
Extensive details are presented in Appendix -H. Here, we briefly present setup, simulation results, and a discussion.
@@ VI-B1 Planning and Tracking Models
As we saw in Section V-A, carefully planning model design is critical to PARC’s reach-avoid. In this case, we fit an affine time-varying model to drifting data collected by rolling out open-loop drifting maneuvers. This not only guarantees a reasonable planning model (i.e., it successfully parameterizes drifting maneuvers) but also demonstrates the flexibility of using PWA models for PARC since they can readily be learned from expert demonstrations.
We use a tracking model based on [jelavic2017autonomous, goh2020toward, weber2023modeling, goh2016simultaneous], detailed in Appendix -H. Our tracking controller is based on [jelavic2017autonomous]: we first use nonlinear MPC to bring the car into a drifting state, then switch to open-loop control to complete the parking action (since the vehicle loses controllability when sliding sideways to a stop).
@@ VI-B3 Results
Figure 12: Example planned drifting trajectories sampled from the Backward Reach-Avoid Set using PARC. The two red cars are obstacles. The goal set for the center of mass is in green. The trajectory with a very low tracking error is shown in (a), whereas a high tracking error is shown in (b). Critically neither collide with the obstacles, and both are represented in the single BRAS.
@@ VI-B4 Discussion
This demo showcased (i) PARC’s capability to generate safe and goal-reaching trajectories with challenging near-danger tasks such as reach-avoid drift parking, and (ii) the flexibility of PARC in terms of handling a suitable planning model from data. We notice that the BRAS is relatively small (see Fig. 1(b)) in the state space, indicating that drift parking is indeed a difficult maneuver. This emphasizes the importance of PARC minimizing conservativeness in the BRAS computation. However, it also suggests that this maneuver is brittle to large changes in initial conditions, and PARC will struggle to find drifting maneuvers for different goal sets or obstacle configurations. In future work, we plan to test more robust drifting controllers to produce a larger BRAS. With that in mind, our ultimate goal is physical hardware experiments on a platform such as the F1:10 car in Fig. 17.

### v4

@@ V-A2 Hypothesis
We expected PARC to be able to compute plans to the goal reliably, with no safety violations during $[0,t_{\mathrm{\textnormal{f}}}]$. However, due to the naïve choice of planning model, we expected the avoid set to be overly conservative, especially since (50) has three non-ETI states ($v_{x},v_{z},\omega$) in its PWA formulation.
@@ V-A4 Discussion
In Appendix -H and the next few examples, we show how a careful design of the planning model leads to much better performance on more complicated scenarios, as well as enabling stabilizing behavior such as braking without explicitly modeling velocity as a goal state.
@@ V-B Impact of Planner-Tracker Cooperation
We now study cooperative vs. adversarial planning and tracking models. We also further study the Neural CLBF [dawson2022safe] learning-based reach-avoid approach. For this experiment, we use a 3-D near-hover quadrotor introduced in FaSTrack [chen2021fastrack], a reachability-based planner tracker that uses adversarial planning and tracking models.
@@ V-B1 Experiment Setup
The tracking model and PARC’s planning model are listed in Appendix -F. The quadrotor must navigate a narrow gap (width 1.8 m, compared to drone width 0.54 m) to reach its goal, as shown in Fig. 7. We simulated 8,100 trials from different initial conditions. We measured the number of trials that reached the goal, and the number that crashed. We also considered the total computation time of each method. Further details are in Appendix -G.
@@ V-B2 Hypotheses
We anticipated that our PARC planner would always reach the goal safely. We also anticipated faster computation time in our reachability analysis compared to the other two approaches, although FaSTrack does not need to recompute the reachable set for every obstacle. Since FaSTrack provides a formal proof of safety, we anticipate it would have no crashes. However, since it treats the planner and tracker as adversarial, we expected it to struggle to reach the goal. On the other hand, since CLBFs specifically balance safety and liveness, we expected it to reach the goal more often than FaSTrack but at the expense of crashing due to the high chance of learning an incorrect representation of safety when forced to operate near obstacles.
@@ V-B3 Results
Fig. 8 illustrates the trajectories simulated on the grid of 8,100 initial positions using each method. Neural CLBF found 576 successful trajectories (7.1%), while safety was violated for the rest. FaSTrack found 0 successful trajectories but never violated safety. However, PARC found the plan to reach the goal without collision for 2,592 initial states (32%) and never violated safety.
@@
Figure 8: Simulated trajectories of 10-D quadrotor for 8,100 initial states using three methods: PARC, FaSTrack [chen2021fastrack], and Neural CLBF [dawson2022safe]. The green box is the goal set and the pink boxes are the obstacles.
@@ V-B4 Discussion
The performance of FaSTrack is notably compromised by the large, conservatively computed TEB, specified as a $2.47\times 2.47\times 0.1$ box. This resulted in its inability to compute feasible plans in spaces tighter than the TEB. However, by sampling tracking error under a parameterized plan, PARC’s tracking error can be much smaller than TEB, as shown in Fig. 9. This enables PARC to find plans even in challenging scenarios.
Although Neural CLBF succeeded in capturing most of the safe, unsafe, and goal regions, the synthesized control policy would violate safety when starting at some of the initial states. Thus, the learning-based method does not provide safety guarantees.
Since PARC has a strong safety guarantee, every plan it found reached the goal and did not collide with the obstacles. We note that a better choice of planning model could increase PARC’s success rate; the simplicity of the single-integrator planning model, employed mainly for fair comparison with FaSTrack, reduced the flexibility of possible plans. In the next section, we will leverage a better choice of trajectory model.
TABLE II: Performance Evaluation of PARC and Other Reach-Avoid Frameworks: This table compares success and safety rates across various systems and scenarios. The success rate is the percentage of simulations achieving their goals without collisions, while the safety rate measures that of obstacle avoidance. Computation times are categorized into offline and online phases. The goals, obstacles, and initial states were known at the beginning of the online phase but not during the offline phase. We report the mean and the standard deviation for computation time.
@@
13-D Quadrotor PARC-Funnel 1.00 1.00 $3.07\pm 0.06$ 00 V-D
@@ V-C1 Experiment Setup
We use the same planning model as quadrotor RTD [kousik2019safe], which generates time-varying polynomial positions in 3-D [mueller2015computationally] with $\Delta t=0.02$units with $t_{\mathrm{\textnormal{f}}}=3$units; this formulation, detailed in Appendix -F2, is affine time-varying in the trajectory parameters with no non-ETI states, so reachability can be tightly approximated by PARC. Both PARC and RTD use the same tracking error that is precomputed in [kousik2019safe].
@@
We declare a trial safe if the quadrotor does not crash, and success if the quadrotor reaches the goal region without colliding with any obstacles.
@@ V-C2 Hypothesis
We anticipated that PARC would safely find ways to reach the goal more often, as its primary source of conservativeness is its tracking error estimation (which is identical to that of RTD). We anticipated that RTD would reach the goal much less often because it suffers an additional source of conservativeness from using a polynomial planning model with zonotopic reachable sets.
@@ V-D Impact of Time-Varying Tracking Error
[…]
Finally, we explore the importance of accounting for the temporal correlation of tracking errors. In this experiment, we compare our approach with the KDF framework that combines sampling-based planning with funnel-based feedback control [verginis2022kdf]. This control mechanism prescribes the tracking error to remain in funnels described by a performance function, $\rho(t)$, which decays exponentially over time [bechlioulis2014low]. Employing the same funnel-based control across both the KDF and PARC methodologies allows us to maintain a consistent tracking error model. This setup confirms that any observed differences in performance are solely attributed to the planner’s strategy of using tracking errors.
Key result: By considering time-varying as opposed to worst-case tracking error, PARC significantly reduces the conservativeness of motion planning. Furthermore, PARC’s use of parameterized trajectories enabled the design of a funnel controller that maintained within the prescribed tracking error bounds, whereas the same controller could not reliably track the smoothed output of a sampling-based motion planner.
@@ V-D1 Experiment Setup
[…]
@@
For the tracker, both PARC and KDF use a specialized funnel controller adapted for an underactuated 3-D quadrotor proposed in [lapandic2022robust], as the conventional KDF funnel-based feedback controller is unsuitable for the system [lapandic2023kinodynamic]. This controller conforms to the performance function $\rho(t)=\rho_{0}+(\rho_{0}-\rho_{\infty})e^{-\lambda t}$ where we choose $\rho_{0}=0.5,\rho_{\infty}=0.1,\lambda=0.5$. We note that this controller does not respect input limits. The 48 hyperparameters for the funnel controllers are reported in Appendix -G3.
Despite extensive tuning efforts, a universal hyperparameter set for KDF could not be determined that respected the prescribed tracking error bounds; thus, we assume the ideal performance of the funnel controller, and base our statistics on reference trajectories ignoring tracking error. In other words, the comparison is biased towards benefiting KDF and disadvantaging PARC. We note that we designed controllers that function for a selected few initial states with KDF, the results of which are shown in Appendix -G3 and Fig. 11. In contrast, since PARC uses parameterized reference trajectories, we were able to succesfully identify a universal set of hyperparameters that enabled us to base the statistics on realized trajectories that included tracking error.
@@ V-D2 Hypothesis
In the wide-gap scenario, we expected both PARC and KDF to successfully find and execute reach-avoid plans. We anticipated KDF to demonstrate faster computation times, benefiting from its sampling-based planning approach. In the narrow-gap scenario, KDF is expected to struggle due to its reliance on maximum funnel values for free space construction in RRT, whereas PARC is expected to be more successful by leveraging the time-varying nature of tracking errors. We expected no safety violations for all experiments.
@@ V-D4 Discussion
As anticipated, PARC demonstrates superior performance over KDF in narrow-gap scenarios due to its ability to utilize the time-varying nature of tracking errors effectively; thus taking advantage of a low tracking-error regime. This advantage is visually depicted in Fig. 11a, where PARC waits for the funnel controller to reduce the tracking error before navigating through narrow gaps. In contrast, for wider gaps shown in Fig. 11b, the quadrotor proceeds through the gap immediately. Conversely, KDF struggles in narrow spaces, as depicted in Fig. 11c since its planner computes paths based on a static maximum tracking error that exceeds the actual gap width (i.e., $\rho_{0}=0.5>0.31$), but performs well in wider spaces where this limitation does not impede planning, as depicted in Fig. 11d.
@@
PARC achieves reach-avoid plans in narrow-gap scenarios but faces lengthy computation times due to detailed sampling from BRAS and smaller $\Delta t$ steps. In contrast, KDF uses sampling-based motion planning effectively in scenarios with adequate gap widths, swiftly creating viable plans. This indicates a potential research direction: integrating sampling-based motion planning within PARC could shorten computation times and facilitate the generation of expert trajectories that adeptly handle near-danger scenarios, surpassing conventional kinodynamic sampling-based motion planning methods such as KDF.
@@ V-E Overall Discussion
Through the above experiments, we have shown how our proposed PARC approach confers advantages in computing and using BRAS. In near-danger scenarios, PARC is less conservative than FaSTrack and KDF since the former uses a time-variant tracking error function, whereas the latter two assume constant, worst-case tracking errors. While RTD can similarly handle time-variant tracking error functions, PARC still outperforms it due to the lower numerical approximation error from representing the planning model as a PWA system instead of a polynomial and using H-polytopes instead of zonotopes. Finally, Neural CLBF cannot provide hard reach-avoid guarantees, resulting in many crashes with obstacles.
[…]
@@ VI-A2 Reach-Avoid Method Comparison
Our drifting tracking model dynamics (see Appendix -I) are not control affine, so directly applying a control barrier function (CBF) or CLBF control approach is not applicable.[ames2016control, dawson2023safe]. Note that [dawson2022safe] poses a control affine vehicle dynamics model but does not include tire saturation, so it cannot drift. Furthermore, our planning and tracking models are too high-dimensional for standard Hamilton-Jacobi-Bellman (HJB) value function computation [chen2021fastrack, mitchell2008flexible]. We also leave a comparison against learning-based HJB [bansal2021deepreach] approaches as future work. It is not obvious to design a valid funnel controller for such an under-actuated, hybrid system as a drifting vehicle to compare with KDF [verginis2022kdf]. The most relevant work is RTD [kousik2020bridging, liu2022refine], which has not been applied to these extreme dynamics. In future work, RTD’s reachable set can use PARC’s formulation to enable a direct drifting comparison, such that RTD’s only difference with PARC is that it implicitly represents the avoid set. Due to this similarity, we focus on demonstrating PARC’s capabilities as opposed to a head-to-head comparison.
@@ VI-B Demo Details
Extensive details are presented in Appendix -I. Here, we briefly present setup, simulation results, and a discussion.
@@ VI-B1 Planning and Tracking Models
As we saw in Section V-A, a careful planning model design is critical to PARC’s reach-avoid guarantees. In this case, we fit an affine time-varying model to drifting data collected by rolling out open-loop drifting maneuvers. This not only guarantees a reasonable planning model (i.e., it successfully parameterizes drifting maneuvers) but also demonstrates the flexibility of using PWA models for PARC since they can readily be learned from expert demonstrations.
We use a tracking model based on [jelavic2017autonomous, goh2020toward, weber2023modeling, goh2016simultaneous], detailed in Appendix -I. Our tracking controller is based on [jelavic2017autonomous]: we first use nonlinear MPC to bring the car into a drifting state, then switch to open-loop control to complete the parking action (since the vehicle loses controllability when sliding sideways to a stop).
@@ VI-B3 Results
Figure 12: Examples of planned drifting trajectories sampled from the Backward Reach-Avoid Set using PARC. The two red cars are the obstacles. The goal set for the center of mass is shown in green. The trajectory with a very low tracking error is shown in (a), whereas that with a high tracking error is shown in (b). Critically, none of the sampled trajectories from the BRAS collide with the obstacles when tracked.
@@ VI-B4 Discussion
This demo showcased (i) PARC’s capability to generate safe and goal-reaching trajectories with challenging near-danger tasks such as reach-avoid drift-parking, and (ii) the flexibility of PARC in being compatible with a planning model learned from data. We notice that the BRAS is relatively small (see Fig. 1(b)) in the state space, indicating that drift parking is indeed a difficult maneuver. This emphasizes the importance for PARC to minimize conservativeness in the BRAS computation. However, it also suggests that this maneuver is brittle to large changes in initial conditions, and PARC will struggle to find drifting maneuvers for different goal sets or obstacle configurations. In future work, we plan to test more robust drifting controllers to produce a larger BRAS. With that in mind, our ultimate goal is physical hardware experiments on a platform such as the F1:10 car in Fig. 17.